# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

### v2

@@ IV Numerical Results
Figure 4: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{max}}=70$ dBm.
Figure 5: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{max}}=70$ dBm.
This section evaluates the performance of the proposed PASS-ISAC framework. A 3D topological network setup is considered, where the dielectric waveguide is located in the x-o-z plane with a height of $d$ and a length of $50$ m. The communicating user and the sensing target are located in a square region centered at the origin in the x-o-y plane. The default simulation parameters are set as: $\sigma^{2}=-105$ dBm, $f=28$ GHz, $d=10$ m, $\Delta x=\frac{\lambda}{2}$, $r_{\text{s}}=30$ m, $\phi_{\text{s}}=\frac{\pi}{3}$, $[x_{1},y_{1},z_{1}]=[-15,-15,0]$ m, $N=16$, $\eta_{\text{eff}}=1.4$, $R_{\text{QoS}}=10$ bps/Hz, $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=10^{-3}$. The other network parameters are shown in the captions of the figures.
@@
  • Conventional antenna: In this scheme, we deploy $N$ conventional uniform linear array (ULA) at the BS, where the antenna space is set as $\frac{\lambda}{2}$.
@@
  • Fixed pinching antenna: In this scheme, $N$ pinching antennas are uniformly spread along the dielectric waveguide.
  • Low-complexity MRC-based scheme: In the low-complexity maximum-ratio combining (MRC)-based scheme, $\mathbf{w}$ is designed by aligning with both $\mathbf{h}_{\text{c}}$ and $\mathbf{h}_{\text{s}}$ simultaneously, i.e., $\mathbf{w}=\sqrt{\frac{P_{\text{max}}}{N}}(\mathbf{h}_{\text{c}}+\mathbf{h}_{\text{s}})\odot\frac{1}{\mathbf{\check{h}}}$, where $\mathbf{\check{h}}=[|\mathbf{h}_{[1],\text{c}}+\mathbf{h}_{[1],\text{s}}|,\cdots,|\mathbf{h}_{[N],\text{c}}+\mathbf{h}_{[N],\text{s}}|]$.
@@
In Fig. 5, we can observe that the pinching antenna achieves the highest illumination power compared to the other baseline schemes. This result can be expected because compared with the conventional antenna and fixed pinching antenna, pinching antennas can be flexibly repositioned to attenuate the large-scale path loss between the pinching antennas and the receiving ends. It indicates that the pinching antenna is capable of providing more spatial degrees-of-freedom (DoFs) to favor the communication and sensing performance. On the other hand, as the low-complexity MRC-based scheme cannot guarantee the optimal radiation pattern design under the given antenna activation positions, it exhibits inferior performance to the proposed scheme.
Fig. 5 depicts the relationship between the illumination power and the number of activated pinching antennas, with a comparison of an upper bound (i.e., each pinching antenna can radiate an arbitrary proportion of the entire transmit power) and proportional power allocation $(P_{n}=\delta(\sqrt{1-\delta^{2}})^{n-1}P_{\text{max}})$ model [10]. Without loss of generality, $\delta$ is set as $0.5$. As can be observed, the equal power allocation model achieves a comparable performance to that of the upper bound. This is because, in the equal power model, each pinching antenna will be allocated a constant level of power, which ensures the spatial DoF gain of the pinching antennas. However, in the proportional power model, pinching antennas located far away from the BS cannot obtain enough power to serve the nearby users, which limits the number of effective pinching antennas. We also observe that increasing $r_{\text{s}}$ decreases the illumination power at the target due to a significant increase in large-scale path loss.
@@
Fig. 5 investigates the impact of the rotation angle of the dielectric waveguide on illumination power at the target. Here, we assume the dielectric waveguide can be rotated in a clockwise direction parallel to the x-o-y plane, where the rotation angle is defined as the angle entwined by the dielectric waveguide and the x-axis. From Fig. 5, it is shown that the illumination power first increases and then decreases as the rotation angle grows. This is due to the fact that when the rotation angle is $60^{\circ}$, the target is located underneath the dielectric waveguide, and it receives the maximum illumination power. As the rotation angle further rises, the distance between the target and the pinching antenna becomes large, so the illumination power gradually decreases. In addition, raising the height of the dielectric waveguide increases the average distance from the pinching antennas to the user and target, thus, the illumination power decreases as $d$ increases.
@@ V Conclusion
A novel PASS-ISAC framework has been proposed, where the pinching beamforming was exploited to realize the simultaneous C&S transmission. A separated ISAC design was proposed for the two-waveguide PASS. A penalty-based AO algorithm was proposed to maximize the illumination power at the target while guaranteeing the QoS requirement of the communication user. Simulation results were provided to verify the superiority of the proposed PASS-ISAC framework over the other benchmarks.

### v3

@@ IV Numerical Results
This section evaluates the performance of the proposed PASS-ISAC framework. A 3D topological network setup is considered, where the dielectric waveguide is located in the x-o-z plane with a height of $d$ and a length of $50$ m. The communicating user and the sensing target are located in a square region centered at the origin in the x-o-y plane. Unless otherwise specified, the default simulation parameters are set as: $\sigma^{2}=-105$ dBm, $f=28$ GHz, $d=10$ m, $r_{\text{s}}=30$ m, $\varphi_{\text{s}}=\frac{\pi}{3}$, $r_{\text{c}}=15\sqrt{2}$ m, $\varphi_{\text{c}}=\frac{5\pi}{4}$, $N=16$, $\eta_{\text{eff}}=1.4$, $R_{\text{QoS}}=10$ bps/Hz, $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=10^{-3}$, and $\alpha_{\text{s}}=1$. The other network parameters are shown in the captions of the figures.
@@
  • Conventional antenna: In this scheme, we deploy $N$ conventional uniform linear array (ULA) at the BS as the transmitting antenna with an antenna spacing of $\frac{\lambda}{2}$. For fairness comparison, the transmitting antennas are connected to one RF chain and each antenna is associated with an analog phase shifter, which can be varied from 00 to $2\pi$.
@@
  • Fixed pinching antenna: In this scheme, $N$ pinching antennas are uniformly spread along the dielectric waveguide, where the in-waveguide and free-space channels are determined by the fixed positions of the pinching antennas.
  • Semi-continuous activation: In the semi-continuous activation scheme, we assume there are $N$ pinching antennas uniformly distributed along the dielectric waveguide, which are predetermined and cannot be changed. However, the pinching antennas are allowed to be adjusted in a small-scale range to alter the phase-shift response of the pinching beamforming, which has a negligible impact on the large-scale path loss.
@@
In Fig. 2, we can observe that the pinching antenna achieves the highest illumination power compared to the other baseline schemes. This result can be expected because, compared with the baseline schemes, pinching antennas can be flexibly repositioned to attenuate the large-scale path loss between the pinching antennas and the receiving ends. Thus, more spatial degrees-of-freedom (DoFs) are provided to favor the communication and sensing performance. On the other hand, although the semi-continuous activation scheme cannot reduce the path loss by adjusting the antenna position over a wide range, it exhibits superior performance to the conventional antenna scheme because pinching antennas are spread over the entire communication/sensing area, which averagely closer to the receiving ends.
Figure 3: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{T}}=70$ dBm.
Fig. 3 depicts the relationship between the illumination power and the number of activated pinching antennas, with a comparison of the proportional power allocation model. For fairness comparison, $\alpha_{\text{s}}=0.9$ for two power allocation models. As can be observed, the illumination power increases as the number of pinching antennas increases, which is because an increasing number of pinching antennas can improve the beam resolution and reduce the power leakage in irrelevant regions, thereby raising the illumination power at the target. It is also observed that the proportional power allocation is slightly inferior to the equal power allocation model, which verifies the effectiveness of the pinching antennas based on proportional power allocation model in reconfiguring signal propagation.
Figure 4: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{T}}=70$ dBm.
@@
Fig. 4 investigates the impact of the rotation angle of the dielectric waveguide on illumination power at the target. Here, we assume the dielectric waveguide can be rotated in a clockwise direction parallel to the x-o-y plane, where the rotation angle is defined as the angle entwined by the dielectric waveguide and the x-axis. From Fig. 4, it is shown that the illumination power first increases and then decreases as the rotation angle grows. This is due to the fact that when the rotation angle is $60^{\circ}$, the target is located underneath the dielectric waveguide, and it receives the maximal illumination power. As the rotation angle further rises, the distance between the target and the pinching antenna becomes large, so the illumination power gradually decreases. In addition, raising the height of the dielectric waveguide increases the average distance from the pinching antennas to the user and target, thus, the illumination power decreases as $d$ increases.
@@ V Conclusion
A novel PASS-ISAC framework has been proposed, where the pinching beamforming was exploited to realize the simultaneous C&S transmission. A separated ISAC design was proposed for the two-waveguide PASS. A penalty-based AO algorithm was proposed to maximize the illumination power at the target while guaranteeing the QoS requirement of the communication user. Simulation results were provided to verify the superiority of the proposed PASS-ISAC framework over the other baseline schemes.